[Brahms-l] Shift report 20050219 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Feb 19 2005 - 23:57:41 EST
Shifter leader was HHD
Shifters were 


1. Good, stable beams were provided by MCR. They even extended the first one for half an hour.

2.  Runs:        13295,  13296, 13306*, 13309  

    FS@4B3450:    205k,  149k, 51k, 207k
    MRS@40A1050:  206k,  260k, 55k, 224k
    begun r13310



3. Problems with Tpc1 DAQ, run 13306 was taken without it.     


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Feb 19 23:57:58 2005

This archive was generated by hypermail 2.1.8 : Sat Feb 19 2005 - 23:58:11 EST